# Chen, Yan, and Jiang

### v1

@@ 1 Introduction
This work is motivated by the pricing challenges faced by such platforms—for example, in setting upfront fares for shared rides, where passengers have different ODs. Pricing and matching are inherently interdependent, naturally forming a bi-level optimization problem: at the upper level, the platform determines optimal prices to maximize profit, often expressed as a concave revenue component minus a total cost component; whereas at the lower level, the platform matches agents to minimize total costs. For a detailed discussion of this application, see Yan, Yan, and Shen (2025).
@@
The lower-level matching problem has received a lot of attention in the literature recently. Existing studies often formulate it as a Markov Decision Process (MDP), under either a continuous-time setting (Aouad and Sarıtaç 2022) or a discrete-time setting (Ashlagi et al. 2023, Gupta 2024, Eom and Toriello 2024, Kerimov, Ashlagi, and Gurvich 2024, 2025). In the MDP framework, the state of the system is defined as the number of agents of each type waiting in the system to be matched. When an agent arrives, the platform can (i) immediately match her with another existing agent, or (ii) keep her waiting in the system. However, the optimal policies are in general intractable to obtain due to the curse of dimensionality. Therefore, many studies use fluid relaxations based on linear programs (LPs) to give bounds for the optimal value function. For example, for the dynamic non-bipartite matching problem, Eom and Toriello (2024), Gupta (2024) and Kerimov, Ashlagi, and Gurvich (2024, 2025) give reward-maximizing LP formulations under infinite agent patience. Collina et al. (2020) and Aouad and Sarıtaç (2022) formulate LPs incorporating limited patience. In this work, we adopt the cost-minimizing LP by Aouad and Sarıtaç (2022) to model $c(\boldsymbol{\lambda})$.
@@
Organization. The remainder of the paper is organized as follows. We introduce the model and preliminaries in Section 2. We perform concave analysis on the value function in 3 and discuss its application on price optimization in 4. We conclude in Section 5. All proofs, auxiliary results, and additional examples are provided in the appendix.
@@ 2 Preliminaries
This LP was originally introduced in Aouad and Sarıtaç (2022), where it was shown that its optimal value provides a lower bound on the expected cost under the optimal matching policy. While Aouad and Sarıtaç (2022) treat $\boldsymbol{\lambda}$ as exogenously given, we instead view this LP as a parametric programming model in $\boldsymbol{\lambda}$. Formally, we consider $c(\boldsymbol{\lambda})$ as a function of $\boldsymbol{\lambda}$ defined over a compact interval $\Lambda=[\underline{\boldsymbol{\lambda}},\overline{\boldsymbol{\lambda}}]$. These lower and upper bounds can be interpreted as minimum and maximum prices imposed on different agent types. For convenience, we assume $\underline{\boldsymbol{\lambda}}>\boldsymbol{0}$, where all vector inequalities in this paper are interpreted component-wise.
@@
When $\theta_{i}=0$, $\forall i\in[N]$, $c(\boldsymbol{\lambda})$ is linear in $\boldsymbol{\lambda}\in\Lambda$.
@@ 3 Concavity of Value Function
(Weak concavity) $c(\boldsymbol{\lambda})$ is said to be weakly concave in $\boldsymbol{\lambda}\in\Lambda$ if there exists a finite value $\rho\geq 0$, such that $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is concave in $\boldsymbol{\lambda}\in\Lambda$.
@@
If $\forall\boldsymbol{\lambda}\in\Lambda$, all optimal solutions satisfy $\boldsymbol{y}^{*}>\boldsymbol{0}$, then $c(\boldsymbol{\lambda})$ is weakly concave in $\boldsymbol{\lambda}\in\Lambda$.
@@ 3.1 Analysis of Two Types of Agents
It is well known that whenever an LP admits an optimal solution, there always exists an optimal basic feasible solution (BFS) (Bertsimas and Tsitsiklis 1997). In any such optimal BFS, a subset of the inequality constraints—specifically, those in (2c), (2d), and (2e)—will be binding. We leverage this fact to characterize the structure of optimal solutions, that is, to determine which constraints are binding for different values of $\lambda_{1}$ and $\lambda_{2}$. For ease of exposition, we define:
@@
Table 1: Characterization of optimal solutions to program (2) when $N=2$
@@
Lemma 3.4 identifies four possible types of optimal solutions, determined by the signs of three indicator functions. When $x_{i,j}^{*}=0$, the corresponding constraint in (2d) is binding; when $x_{i,j}^{\ast}=\lambda_{j}y_{i}^{\ast}/\theta_{i}$, the corresponding constraint in (2c) is binding. The first solution category corresponds to the case with no cross-matching between type 1 and type 2 agents (i.e., $x_{1,2}^{\ast}=x_{2,1}^{\ast}=0$). The second category is the fully-matched case, where all types of matchings occur and all constraints in (2c) are binding. In the third category, there is no self-matching among type 1 agents ($x_{1,1}^{\ast}=0$). The last category is characterized by $y_{2}^{\ast}=0$, where all type 2 agents are passively matched with type 1 agents. Using Lemma 3.4 and flow constraints (2b), we can explicitly give the expression of $y_{1}^{*}$ and $y_{2}^{*}$, thereby obtaining the optimal solutions in closed forms. The complete closed-form expressions are provided at the end of Appendix 6.4. Lemma 3.4 also helps us to identify the region where $c(\boldsymbol{\lambda})$ is weakly concave. Based on the results of Lemma 3.3 and Lemma 3.4, we see that if, for every $(\lambda_{1},\lambda_{2})\in\Lambda$, at least one of the inequalities $\Delta_{2}(\lambda_{1},\lambda_{2})<0$ or $\Delta_{3}(\lambda_{1},\lambda_{2})<0$ holds, then $y_{1}^{\ast},y_{2}^{\ast}>0$ is guaranteed. The following theorem further explores this condition and gives an explicit characterization of $\Lambda$ satisfying such conditions.
@@
$\tau_{2}$ $:=\tau_{1}^{2}-8c_{(1)}(2c_{(1,2)}-c_{(1)}-c_{(2)})\theta_{1}(\theta_{1}+\theta_{2}).$
(i) When $\tau_{1}\leq 0$ or $\tau_{2}<0$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$;
(ii) When $\tau_{1}>0$ and $\tau_{2}\geq 0$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$ if $2c_{(1,2)}-c_{(1)}-c_{(2)}\neq 0$ and
$\underline{\lambda}_{2}>\frac{\tau_{1}+\sqrt{\tau_{2}}}{4(2c_{(1,2)}-c_{(1)}-c_{(2)})}.$
Theorem 3.5 gives two sufficient conditions for weak concavity, depending on the signs of $\tau_{1}$ and $\tau_{2}$. Smaller values of $\tau_{1}$ and $\tau_{2}$ are desired to achieve weak concavity. The expressions of $\tau_{1}$ and $\tau_{2}$ lead to the following insights: (i) Since both $\tau_{1}$ and $\tau_{2}$ are decreasing in $\theta_{1}$, a larger value of $\theta_{1}$ (i.e., when $\theta_{1}$ and $\theta_{2}$ are closer since $\theta_{1}\leq\theta_{2}$) helps to achieve weak concavity in the larger region; (ii) Similarly, since both $\tau_{1}$ and $\tau_{2}$ are decreasing in $c_{(1,2)}$, a larger value of $c_{(1,2)}$ also promote weak concavity. Although Theorem 3.5 provides only sufficient conditions, we find that the resulting lower bound for $\underline{\lambda}_{2}$ is quite tight. In Appendix 7, we present a counterexample (Example 7.3) demonstrating that as soon as $\lambda_{2}$ falls just below our derived bound, weak concavity is already violated.
Note that in the second case of Theorem 3.5, we require $2c_{(1,2)}-c_{(1)}-c_{(2)}\neq 0$, or equivalently, $e_{1,2}\neq 0.5$. This condition means that the two agent types cannot have perfect matching efficiency. In the context of ridesharing, this essentially precludes the possibility that both rider types have exactly the same OD pairs. When $e_{1,2}=0.5$ and $\theta_{1}=\theta_{2}$, this is not an issue, as the platform can simply merge the two types into a single type. However, as the following proposition shows, if $\theta_{1}\neq\theta_{2}$, the region where $y_{2}^{*}=0$ becomes unbounded, making it impossible to establish any lower bound on $\underline{\boldsymbol{\lambda}}$ that guarantees $\boldsymbol{y}^{*}>\boldsymbol{0}$. Appendix 7 provides a counterexample (Example 7.4) demonstrating that weak concavity can easily be violated when $e_{1,2}=0.5$ and $\theta_{1}\neq\theta_{2}$, even for large values of $\lambda_{1}$ and $\lambda_{2}$.
@@
(i) When $\theta_{2}<3\theta_{1}$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$;
(ii) When $\theta_{2}\geq 3\theta_{1}$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$ if $e_{1,2}\neq 0.5$ and
$\underline{\lambda}_{2}>\frac{\theta_{2}}{4}\frac{1}{1-2e_{1,2}}.$
According to Corollary 3.7, the function is guaranteed to be weakly concave as long as $\theta_{2}<3\theta_{1}$. If the two agent types have significantly different patience levels, we require a lower bound on $\underline{\lambda}_{2}$. This bound depends on $\theta_{2}$ (the patience of agent type 2) and $e_{1,2}$ (the matching efficiency between the two agent types). A smaller $\theta_{2}$ (greater patience for both types of agents since $\theta_{1}\leq\theta_{2}$) and a lower $e_{1,2}$ (lower matching efficiency) both expand the region of $\Lambda$ where weak concavity is maintained.
@@
Table 2: Characterization of optimal solutions to program (2) when $N=2$ and $\theta_{1}=\theta_{2}=\theta$
@@
When $N=2$ and $\theta_{1}=\theta_{2}=\theta$, $c(\boldsymbol{\lambda})$ is concave in $\boldsymbol{\lambda}\in\Lambda$.
@@ 3.2 Analysis of Multiple Types of Agents
We now present results for the general case with an arbitrary number of agent types. As $N$ increases, fully characterizing the optimal solutions becomes intractable. However, by leveraging insights from previous results and their proofs, we can identify sufficient conditions for ensuring $\boldsymbol{y}^{*}>\boldsymbol{0}$ for general $N$, and once again apply Lemma 3.3 to establish weak concavity. For clarity, we assume throughout this section that $\theta_{1}\leq\theta_{2}\leq\cdots\leq\theta_{N}$.
Given any type $i$, we first sort its corresponding matching efficiency $e_{i,1},e_{i,2},\cdots,e_{i,N}$ in descending order and denote the sorted result as $e_{i,(1)}\geq e_{i,(2)}\geq\cdots\geq e_{i,(N)}$, where by definition $e_{i,(1)}=e_{i,i}=0.5$. For $k\in\mathbb{Z}^{+}$, we define the $k$-th critical matching efficiency as
$e_{(k)}=\begin{cases}\max\limits_{i\in[N]}\left\{e_{i,(k)}\right\},&\text{if }k\leq N,\\ 0,&\text{if }k>N,\end{cases}$
that is, when $k\leq N$, $e_{(k)}$ is the maximum $k$-th largest matching efficiency among all agent types. Since each agent has a self-matching efficiency of $0.5$, we always have $e_{(1)}=0.5$. When $k\geq 2$, this implies that for any agent of type $i\in[N]$, there are at most $k-2$ other agent types (excluding self-match type $i$) with a matching efficiency greater than $e_{(k)}$. For notational convenience, we also define $e_{(k)}=0$ for all $k>N$.
@@
(i) When $\theta_{1}=\theta_{2}=\cdots=\theta_{N}=\theta$, $c(\boldsymbol{\lambda})$ is concave in $\Lambda$ if $e_{(3)}\neq 0.5$ and
$\underline{\lambda}_{i}>\theta\frac{e_{(3)}}{1-2e_{(3)}},\ \forall i\in[N];$
(ii) When $\theta_{1}\neq\theta_{N}$, $c(\boldsymbol{\lambda})$ is concave in $\Lambda$ if $e_{(2)}\neq 0.5$ and
$\underline{\lambda}_{i}>\theta_{i}\frac{e_{(2)}}{1-2e_{(2)}},\ \forall i\in[N].$
@@
(i) When $\theta_{1}=\theta_{2}=\cdots=\theta_{N}=\theta$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$ if $e_{(4)}\neq 0.5$ and
$\underline{\lambda}_{i}>\theta\frac{e_{(4)}}{1-2e_{(4)}},\ \forall i\in[N];$
(ii) When $\theta_{N}<2\theta_{1}$, $c(\boldsymbol{\lambda})$ is weakly concave in $\Lambda$ if $e_{(3)}\neq 0.5$ and
$\underline{\lambda}_{i}>\theta_{i}\frac{e_{(3)}}{1-2e_{(3)}},\ \forall i\in[N].$
Similar to Theorem 3.10, we also require a non-zero lower bound of $\underline{\lambda}_{i}$ depending on the critical matching efficiency $e_{(3)}$ or $e_{(4)}$ in Theorem 3.11. For the first case, we continue to assume that $\theta_{i},\ i\in[N]$ are all equal, but allow the region $\Lambda$ being larger than that in Theorem 3.10. Specifically, the lower bound of $\underline{\lambda}_{i}$ is now related to $e_{(4)}$ instead of $e_{(3)}$. Under this weaker assumption, we can prove that the value function is weakly concave, if not concave. In the second case, we assume the values of $\theta_{i},\ i\in[N]$ can be different but within a ratio of 2, and also gives a condition of weak concavity. Notably, the allowable region is also enlarged compared to Theorem 3.10, since the lower bound for $\underline{\lambda}_{i}$ is determined by $e_{(3)}$ instead of $e_{(2)}$. Both results generalize the key steps used in the proof of Lemma 3.4.
From the first case of Theorem 3.11, we immediately get the following corollary for the case of $N=3$, since $e_{(4)}=0$ according to our definition.
@@
When $N=3$ and $\theta_{1}=\theta_{2}=\theta_{3}$, $c(\boldsymbol{\lambda})$ is weakly concave in $\boldsymbol{\lambda}\in\Lambda$.
Both Theorem 3.10 and Theorem 3.11 state that it is easier to achieve (weak) concavity when all agents have the same patience level. A higher patience level also leads to a smaller required lower bound of $\underline{\lambda}_{i}$ for each $i\in[N]$. As an extreme case, when $\theta_{i}=0$, $\forall i\in[N]$, the value function will become linear, as we previously established in Proposition 2.2.

### v2

@@ 1 Introduction
This work is motivated by the pricing challenges faced by such platforms—for example, in setting upfront fares for shared rides, where passengers have different ODs. Pricing and matching are inherently interdependent (see, e.g., Feng, Niazadeh, and Saberi 2024), naturally forming a bi-level optimization problem: at the upper level, the platform determines optimal prices to maximize profit, often expressed as a concave revenue component minus a total cost component; whereas at the lower level, the platform matches agents to minimize total costs. For a detailed discussion of this shared rides pricing application, see Yan, Yan, and Shen (2025).
@@
The lower-level matching problem has received a lot of attention in the literature recently. Existing studies often formulate it as a Markov Decision Process (MDP), under either a continuous-time setting (Ashlagi et al. 2019, Aouad and Sarıtaç 2022) or a discrete-time setting (Ashlagi et al. 2023, Gupta 2024, Eom and Toriello 2024, Kerimov, Ashlagi, and Gurvich 2024, 2025). In the MDP framework, the state of the system is defined as the number of agents of each type waiting in the system to be matched. When an agent arrives, the platform can (i) immediately match her with another existing agent, or (ii) keep her waiting in the system. However, the optimal policies are in general intractable to obtain due to the curse of dimensionality. Therefore, many studies use fluid relaxations based on linear programs (LPs) to give bounds for the optimal value function. For example, for the dynamic non-bipartite matching problem, Eom and Toriello (2024), Gupta (2024) and Kerimov, Ashlagi, and Gurvich (2024, 2025) give reward-maximizing LP formulations under infinite agent patience. Collina et al. (2020) and Aouad and Sarıtaç (2022) formulate LPs incorporating limited patience. In this work, we adopt the cost-minimizing LP by Aouad and Sarıtaç (2022) to model $c(\boldsymbol{\lambda})$.
@@
Organization. The remainder of the paper is organized as follows. We introduce the model and preliminaries in Section 2. We perform concave analysis on the value function in Section 3 and discuss its application on price optimization in Section 4. We conclude in Section 5. All proofs, auxiliary results, and additional examples are provided in the appendix.
@@ 2 Preliminaries
This LP was originally introduced in Aouad and Sarıtaç (2022), where it was shown that its optimal value provides a lower bound on the expected cost under the optimal matching policy. Correspondingly, using this LP, the optimal profit given in (1) provides an upper bound for the profit achievable by any dynamic matching policy. While Aouad and Sarıtaç (2022) treat $\boldsymbol{\lambda}$ as exogenously given, we instead view this LP as a parametric programming model in $\boldsymbol{\lambda}$. Formally, we consider $c(\boldsymbol{\lambda})$ as a function of $\boldsymbol{\lambda}$ defined over a compact interval $\Lambda=[\underline{\boldsymbol{\lambda}},\overline{\boldsymbol{\lambda}}]$. These lower and upper bounds can be interpreted as minimum and maximum prices imposed on different agent types. For convenience, we assume $\underline{\boldsymbol{\lambda}}>\boldsymbol{0}$, where all vector inequalities in this paper are interpreted component-wise.
@@
When $\theta_{i}=0$, $\forall i\in[N]$, $c(\boldsymbol{\lambda})$ is linear on $\Lambda$.
@@ 3 Concavity of Value Function
(Weak concavity) $c(\boldsymbol{\lambda})$ is said to be weakly concave on $\Lambda$ if there exists a finite value $\rho\geq 0$, such that $c(\boldsymbol{\lambda})-\rho\|\boldsymbol{\lambda}\|^{2}/2$ is concave on $\Lambda$.
@@
If $\forall\boldsymbol{\lambda}\in\Lambda$, all optimal solutions satisfy $\boldsymbol{y}^{*}>\boldsymbol{0}$, then $c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$.
@@ 3.1 Analysis of Two Types of Agents
It is well known that whenever an LP admits an optimal solution, there always exists an optimal BFS (Bertsimas and Tsitsiklis 1997). In any such optimal BFS, a subset of the inequality constraints—specifically, those in (2c), (2d), and (2e)—will be binding. We leverage this fact to characterize the structure of optimal solutions, that is, to determine which constraints are binding for different values of $\lambda_{1}$ and $\lambda_{2}$. For ease of exposition, we define:
@@
Table 1: Characterization of optimal solutions to program (2) when $N=2$.
@@
Lemma 3.4 identifies four possible types of optimal solutions, determined by the signs of three indicator functions. When $x_{i,j}^{*}=0$, the corresponding constraint in (2d) is binding; when $x_{i,j}^{\ast}=\lambda_{j}y_{i}^{\ast}/\theta_{i}$, the corresponding constraint in (2c) is binding. The first solution category corresponds to the case with no cross-matching between type 1 and type 2 agents (i.e., $x_{1,2}^{\ast}=x_{2,1}^{\ast}=0$). The second category is the fully-matched case, where all types of matchings occur and all constraints in (2c) are binding. In the third category, there is no self-matching among type 1 agents ($x_{1,1}^{\ast}=0$). The last category is characterized by $y_{2}^{\ast}=0$, where all type 2 agents (with low patience) are passively matched with type 1 agents (with high patience). Using Lemma 3.4 and flow constraints (2b), we can explicitly give the expression of $y_{1}^{*}$ and $y_{2}^{*}$, thereby obtaining the optimal solutions in closed forms. The complete closed-form expressions are provided at the end of Appendix 6.4. Lemma 3.4 also helps us to identify the region where $c(\boldsymbol{\lambda})$ is weakly concave. Based on the results of Lemma 3.3 and Lemma 3.4, we see that if, for every $(\lambda_{1},\lambda_{2})\in\Lambda$, at least one of the inequalities $\Delta_{2}(\lambda_{1},\lambda_{2})<0$ or $\Delta_{3}(\lambda_{1},\lambda_{2})<0$ holds, then $y_{1}^{\ast},y_{2}^{\ast}>0$ is guaranteed. The following theorem further explores this condition and gives an explicit characterization of $\Lambda$ satisfying such conditions.
@@
$\tau_{2}$ $:=\tau_{1}^{2}-8c_{(1)}(2c_{(1,2)}-c_{(1)}-c_{(2)})\theta_{1}(\theta_{1}+\theta_{2}),$
then $c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$ under any of the following conditions:
(i) $\tau_{1}\leq 0$ or $\tau_{2}<0$;
(ii) $\tau_{1}>0$, $\tau_{2}\geq 0$, $2c_{(1,2)}-c_{(1)}-c_{(2)}\neq 0$ and $\underline{\lambda}_{2}>(\tau_{1}+\sqrt{\tau_{2}})/\left[4(2c_{(1,2)}-c_{(1)}-c_{(2)})\right];$
(iii) $\tau_{1}>0$, $\tau_{2}\geq 0$, and $\underline{\lambda}_{2}>\overline{\lambda}_{1}-\theta_{1}$.
Theorem 3.5 gives three sufficient conditions for weak concavity, depending on the signs of $\tau_{1}$ and $\tau_{2}$. For the first case when $\tau_{1}\leq 0$ or $\tau_{2}<0$, $c(\boldsymbol{\lambda})$ is weakly concave on any closed interval $\Lambda$. For the second case, we require a lower bound of $\underline{\lambda}_{2}$, and for the third case, we pose restrictions on both $\underline{\lambda}_{2}$ and $\overline{\lambda}_{1}$. In general, smaller values of $\tau_{1}$ and $\tau_{2}$ are desired to achieve weak concavity. Since both $\tau_{1}$ and $\tau_{2}$ are decreasing in $\theta_{1}$, a larger value of $\theta_{1}$ (i.e., when $\theta_{1}$ and $\theta_{2}$ are closer since $\theta_{1}\leq\theta_{2}$) helps to achieve weak concavity in the larger region. Similarly, since both $\tau_{1}$ and $\tau_{2}$ are decreasing in $c_{(1,2)}$, a larger value of $c_{(1,2)}$ (i.e., lower matching efficiency) also promotes weak concavity. Although Theorem 3.5 provides only sufficient conditions, we find that the resulting bounds are quite tight. In Appendix 7, we also present a counterexample (Example 7.3) demonstrating that as soon as $\lambda_{2}$ falls just below our derived bound in the second case, weak concavity is already violated.
Also note that in the second case of Theorem 3.5, we require $2c_{(1,2)}-c_{(1)}-c_{(2)}\neq 0$, or equivalently, $e_{1,2}\neq 0.5$. This condition means that the two agent types cannot have perfect matching efficiency. In the context of ridesharing, this essentially precludes the possibility that both rider types have exactly the same OD pairs. When $e_{1,2}=0.5$ and $\theta_{1}=\theta_{2}$, this is not an issue, as the platform can simply merge the two types into a single type. However, as the following proposition shows, if $\theta_{1}\neq\theta_{2}$, the region where $y_{2}^{*}=0$ becomes unbounded, making it impossible to establish any lower bound on $\underline{\boldsymbol{\lambda}}$ that guarantees $\boldsymbol{y}^{*}>\boldsymbol{0}$, without restricting $\overline{\boldsymbol{\lambda}}$. In Appendix 7, we also provide a counterexample (Example 7.4) demonstrating that weak concavity can be violated when $e_{1,2}=0.5$ and $\theta_{1}\neq\theta_{2}$, even for large values of $\lambda_{1}$ and $\lambda_{2}$.
@@
$c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$ under any of the following conditions:
(i) $\theta_{2}/\theta_{1}<3$;
(ii) $\theta_{2}/\theta_{1}\geq 3$, $e_{1,2}\neq 0.5$ and $\underline{\lambda}_{2}>\theta_{2}/\left[4(1-2e_{1,2})\right]$;
(iii) $\theta_{2}/\theta_{1}\geq 3$, and $\overline{\lambda}_{1}\leq\theta_{1}$.
According to Corollary 3.7, the function is guaranteed to be weakly concave as long as $\theta_{2}<3\theta_{1}$. If the two agent types have significantly different patience levels, we require a lower bound on $\underline{\lambda}_{2}$ or an upper bound on $\overline{\lambda}_{1}$. The lower bound of $\underline{\lambda}_{2}$ depends on $\theta_{2}$ (the patience of agent type 2) and $e_{1,2}$ (the matching efficiency between the two agent types). A smaller $\theta_{2}$ (greater patience for both types of agents since $\theta_{1}\leq\theta_{2}$) and a lower $e_{1,2}$ (lower matching efficiency) both expand the region of $\Lambda$ where weak concavity is maintained. The upper bound of $\overline{\lambda}_{1}$ depends on $\theta_{1}$, the patience of agent type 1, and when $\theta_{1}$ is larger (i.e., when $\theta_{1}$ and $\theta_{2}$ are closer since $\theta_{1}\leq\theta_{2}$), weak concavity is easier to achieve.
@@
Table 2: Characterization of optimal solutions to program (2) when $N=2$ and $\theta_{1}=\theta_{2}=\theta$.
@@
When $N=2$ and $\theta_{1}=\theta_{2}=\theta$, $c(\boldsymbol{\lambda})$ is concave on $\Lambda$.
@@ 3.2 Analysis of Multiple Types of Agents
We now present results for the general case with an arbitrary number of agent types. As $N$ increases, fully characterizing the optimal solutions becomes intractable. However, by leveraging insights from previous results and their proofs, we can identify sufficient conditions for ensuring $\boldsymbol{y}^{*}>\boldsymbol{0}$ for general $N$, and once again apply Lemma 3.3 to establish weak concavity. For clarity, we assume throughout this subsection that $\theta_{1}\leq\theta_{2}\leq\cdots\leq\theta_{N}$. We also define $\nu:=\theta_{N}/\theta_{1}\geq 1$, that is, a larger $\nu$ represents a larger difference in agents’ patience levels.
Given any type $i$, we sort its corresponding matching efficiency $e_{i,1},e_{i,2},\cdots,e_{i,N}$ in descending order and denote the sorted result as $e_{i,(1)}\geq e_{i,(2)}\geq\cdots\geq e_{i,(N)}$, where by definition $e_{i,(1)}=e_{i,i}=0.5$. For $k\in\mathbb{Z}^{+}$, we define the $k$-th critical matching efficiency as
$e_{(k)}:=\begin{cases}\max\limits_{i\in[N]}\left\{e_{i,(k)}\right\},&\text{if }k\leq N,\\ 0,&\text{if }k>N,\end{cases}$
that is, when $k\leq N$, $e_{(k)}$ is the maximum $k$-th largest matching efficiency among all agent types. Since each agent has a self-matching efficiency of $0.5$, we always have $e_{(1)}=0.5$. When $k\geq 2$, this implies that for any agent of type $i\in[N]$, there are at most $k-2$ other agent types (excluding self-match type $i$) with a matching efficiency greater than $e_{(k)}$. For notational convenience, we also define $e_{(k)}=0$ for all $k>N$. It can be easily verified that $e_{(k)}$ is non-increasing in $k\in\mathbb{Z}^{+}$.
@@
$c(\boldsymbol{\lambda})$ is concave on $\Lambda$ under either of the following conditions:
(i) $\nu=1$, $e_{(3)}\neq 0.5$ and $\underline{\lambda}_{i}>\theta_{i}e_{(3)}/(1-2e_{(3)}),\ \forall i\in[N];$
(ii) $\nu>1$, $e_{(2)}\neq 0.5$ and $\underline{\lambda}_{i}>\theta_{i}e_{(2)}/(1-2e_{(2)}),\ \forall i\in[N].$
@@
$c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$ under either of the following conditions:
(i) there exists $K\in\mathbb{Z}^{+}$, such that $K\nu\leq 2$, $e_{(K+2)}\neq 0.5$, and
$\underline{\lambda}_{i}>\theta_{i}\frac{e_{(K+2)}}{1-2e_{(K+2)}},\ \forall i\in[N];$
(ii) there exists $K\in\mathbb{Z}^{+}$, such that $K\nu>2$, $e_{(K+2)}\neq 0.5$ and
$\theta_{i}\frac{e_{(K+2)}}{1-2e_{(K+2)}}<\underline{\lambda}_{i}<\overline{\lambda}_{i}<\theta_{i}\max\left\{\frac{1}{K},\frac{1}{K\nu-2}\right\}\left(1+\frac{1}{\nu}\frac{e_{(K+2)}}{1-2e_{(K+2)}}\right),\ \forall i\in[N].$
Theorem 3.11 provides two sufficient conditions of weak concavity, depending on the choice of $K$ and the value of $\nu$. The parameter $K$ determines the maximum number of other agent types that can be matched with agent type $i\in[N]$, and $\nu$ measures the difference in patience levels across agent types. When $K\nu\leq 2$, it suffices to impose only lower bounds $\underline{\lambda}_{i}$, $i\in[N]$, whereas when $K\nu>2$, we pose both lower and upper bounds. As $K$ increases, $e_{(K+2)}$ decreases, leading to lower values for the bounds—thus shifting the feasible region $\Lambda$. The value of $K$ can therefore be selected by comparing these feasible regions and identifying the one that is most practically useful. On the other hand, as $\nu$ decreases—indicating more similar patience levels among agents—the upper bounds (when $K\nu>2$) become less restrictive, expanding the allowable region. Therefore, a $\nu$ close to 1 helps to achieve weak concavity.
From Theorem 3.11, we can get the following corollary, by letting $K=2$ when $\nu=1$ and letting $K=1$ when $21<\nu<$ 2.
@@
$c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$ under either of the following conditions:
(i) $\nu=1$, $e_{(4)}\neq 0.5$ and $\underline{\lambda}_{i}>\theta_{i}e_{(4)}/(1-2e_{(4)}),\ \forall i\in[N];$
(ii) $21<\nu<$ 2, $e_{(3)}\neq 0.5$ and $\underline{\lambda}_{i}>\theta_{i}e_{(3)}/(1-2e_{(3)}),\ \forall i\in[N].$
Interestingly, Corollary 3.12 closely resembles Theorem 3.10. For the first case, we continue to assume that $\theta_{i},\ i\in[N]$ are all equal, but allow the region $\Lambda$ being larger than that in Theorem 3.10. Specifically, the lower bound of $\underline{\lambda}_{i}$ is now related to $e_{(4)}$ instead of $e_{(3)}$. Under this weaker assumption, we can prove that the value function is weakly concave, if not concave. In the second case, we assume the values of $\theta_{i},\ i\in[N]$ can be different but within a ratio of 2, and also give a condition of weak concavity. Notably, the allowable region is also enlarged compared to Theorem 3.10, since the lower bound for $\underline{\lambda}_{i}$ is determined by $e_{(3)}$ instead of $e_{(2)}$.
From the first case of Corollary 3.12, we can also get the following corollary for the case of $N=3$, since $e_{(4)}=0$ according to our definition. Therefore, weak concavity always holds when there are fewer than or equal to three agent types with identical patience levels.
Corollary 3.13
When $N=3$ and $\theta_{1}=\theta_{2}=\theta_{3}$, $c(\boldsymbol{\lambda})$ is weakly concave on $\Lambda$.
To summarize, Theorems 3.10 and 3.11, along with Corollary 3.12, all suggest that it is easier to achieve (weak) concavity when agents have similar or identical patience levels (i.e., when $\nu$ is close to or equal to $1$). Additionally, a higher overall patience level or a greater heterogeneity in matching efficiencies also leads to a smaller required lower bound of $\underline{\lambda}_{i}$ for each $i\in[N]$. These findings are consistent with the results established in Section 3.1 when $N=2$. Lastly, as an extreme case, when $\theta_{i}=0$, $\forall i\in[N]$, the value function will become linear, as we previously established in Proposition 2.2.